# Hardware-Assisted Virtualization of Neural Processing Units for Cloud Platforms

### v1

@@ Abstract
We present TCloud, a holistic NPU virtualization framework. We investigate virtualization techniques for NPUs across the entire software and hardware stack. TCloud consists of (1) a flexible NPU abstraction called vNPU, which enables fine-grained virtualization of the heterogeneous compute units in a physical NPU (pNPU); (2) a vNPU resource allocator that enables pay-as-you-go computing model and flexible vNPU-to-pNPU mappings for improved resource utilization and cost-effectiveness; (3) an ISA extension of modern NPU architecture for facilitating fine-grained tensor operator scheduling for multiple vNPUs. We implement TCloud based on a production-level NPU simulator. Our experiments show that TCloud improves the throughput of ML inference services by up to 1.4$\times$ and reduces the tail latency by up to 4.6$\times$, while improving the NPU utilization by 1.2$\times$ on average, compared to state-of-the-art NPU sharing approaches.
@@ I Introduction
Ideally, we wish to virtualize NPUs to enable flexible and fine-grained resource sharing and scheduling for improved NPU utilization and performance isolation. We present TCloud, a hardware-assisted system virtualization framework for NPUs.
@@
TCloud can map vNPUs to physical compute units of NPU cores in different manners, based on the service level objectives (SLOs) of ML services. To maximize the NPU utilization while ensuring performance isolation, TCloud enables fine-grained spatial sharing with resource harvesting. It also enables the oversubscription of NPU cores by temporally sharing MEs/VEs among multiple vNPUs. Therefore, the idle compute units can be opportunistically utilized by collocated workloads.
To facilitate the dynamic scheduling for collocated vNPUs, TCloud extends the VLIW-style ISA by reorganizing VLIW instructions into independent micro-Tensor operators ($\mu$TOps in §$\S$§III). TCloud introduces necessary architectural logic for fine-grained dynamic scheduling of $\mu$TOps on the shared physical NPU cores. It allows one vNPU to harvest available compute cycles of MEs/VEs from collocated vNPUs, without causing much interference. This is impossible with conventional VLIW-style ISAs, as they strictly couple the control flows of the (statically) allocated compute units. Our new architectural support enables TCloud to offer the flexibility of NPU resource allocation and scheduling across the software (i.e., vNPU abstraction) and hardware (i.e., fine-grained $\mu$TOp scheduling) stack. TCloud requires minimum modifications to NPU chips (0.04% die area cost) as well as ML compilers.
@@
We implement TCloud with a production-level NPU simulator following the typical TPU architecture. We collect the traces of ML services as we run the MLPerf benchmarks [48] and the TPU reference models [23] on the real Google TPUs. Our experiments with multi-tenant ML instances show that TCloud can improve the throughput of ML inference services by up to 1.4$\times$ and reduce the tail latency by up to 4.6$\times$, while improving the NPU utilization by 1.2$\times$ on average, in comparison with state-of-the-art NPU sharing approaches. We summarize the contributions of TCloud as follows:
@@ III Design and Implementation
We design TCloud to achieve the following objectives:
@@
vNPU abstraction. A vNPU instance reflects the hierarchy of a physical NPU board. Figure 10 shows the configurable parameters of a vNPU. Each vNPU is exposed to the VM as a PCIe device. The guest NPU driver can query the hierarchy of the vNPU, such as the number of chips, cores per chip, HBM size, and others. The maximum vNPU size is capped by the physical NPU size. If a guest VM requires more resources than is available on a physical NPU board, TCloud can allocate multiple vNPU instances to it. The guest ML framework can handle the data distribution across multiple vNPU cores in the same way as that on physical NPUs. Take Google TPU for example, TensorFlow already handles data parallelism across physical NPUs. It can work in the same way with vNPUs.
Figure 11: System architecture of TCloud.
vNPU lifecycle. To create a vNPU instance, a user can specify the vNPU configuration following the pay-as-you-go model [50]. Cloud providers can define various default configurations (e.g., small/medium/large vNPU cores as having 1/4/8 MEs/VEs). TCloud can also learn an optimized vNPU configuration for a DNN workload with ML compilers (§$\S$§III-B). As shown in Figure 11, upon vNPU initialization, the guest driver sends a request to the hypervisor through a para-virtualized interface (§III-F) ( 1). The vNPU manager maps the vNPU instance to NPU hardware resources (§$\S$§III-C). Then, it initializes the vNPU context in the physical NPU device and creates the MMIO mappings for the guest VM to access the vNPU ( 2). During execution, the application issues commands such as memcpy and compute offloading through the command buffer. The NPU hardware directly fetches the commands from the host memory without the hypervisor intervention. It also has DMA access to the DMA buffer in the guest memory space via the IOMMU. The DNN program on the NPU executes asynchronously from the CPU program, and the NPU hardware schedules vNPUs (§$\S$§III-E) independently of existing OS/hypervisor schedulers. The guest VM waits for the completion interrupt or actively polls the memory-mapped control registers for the current status of the vNPU ( 3). After execution, the user can free the vNPU.
@@
Following the popular pay-as-you-go model [50], cloud platforms allow users to specify the vNPU configuration on demand. However, as ML inference workloads have diverse ME/VE demands (see §$\S$§II-B), specifying the number of MEs/VEs can be challenging for users who are not NPU experts. Thus, we allow them to specify the total number of execution units (EUs), which is directly related to the cost of running the vNPU instance. TCloud provides the vNPU allocator, a compile-time tool to improve the performance per cost of vNPUs by identifying an optimized ME/VE ratio for the user workload.
@@
Memory allocation. Users can use the compiler to estimate the total HBM capacity needed by a DNN workload. By default, the SRAM capacity is allocated proportionally to the number of allocated MEs, as more MEs usually indicate larger tile sizes. Based on our study in §$\S$§II-B, for many common ML inference services, the HBM bandwidth is less of a concern. Thus, TCloud allows fair sharing of HBM bandwidth by default. For large models that demand large HBM capacity and bandwidth, the vNPU abstraction offers the flexibility for end users to allocate the demanded resources. The user may also leverage existing tensor swapping techniques to support large DNN workloads with limited memory capacity [28, 63]. After vNPU allocation, ML compilers will compile the DNN program with the allocated resources. The compiler ensures the DNN program does not exceed the allocated SRAM and HBM. We will discuss how TCloud handles compilation for different numbers of MEs/VEs in §$\S$§III-D.
@@
The vNPU manager attempts to balance the number of allocated EUs and the size of allocated memory. This minimizes the chance that all EUs on one core are allocated but a large portion of its memory is not allocated, or vice versa. Thus, vNPUs with many EUs and small memory will be collocated with vNPUs with few EUs and large memory. TCloud uses a greedy algorithm for this by default.
vNPU mapping schemes. TCloud provides the flexibility for cloud platforms to enable both hardware-isolated (spatial-isolated) mapping and software-isolated (temporal-sharing) mapping. With hardware-isolated mapping, a vNPU is mapped to dedicated EUs and SRAM, and the allocated hardware is not shared with other vNPUs. With software-isolated mapping, multiple vNPUs can temporally share the same EUs. TCloud uses priority-based scheduling for fair sharing and performs context switches between vNPUs (see §$\S$§III-E).
vNPU mapping policies. TCloud decides which vNPUs can be mapped onto the same physical NPU (pNPU) as follows. With hardware-isolated mapping, TCloud collocates a set of vNPUs as long as the total resource requirement (e.g., number of MEs/VEs, HBM capacity) does not exceed the pNPU. With software-isolated mapping, TCloud aims to load-balance the pNPUs while allowing oversubscription. TCloud tracks the total resource requirement of assigned vNPUs on each pNPU, and assigns a new vNPU to the pNPU that suffers the least resource requirement. TCloud can support other collocation policies [41, 57, 12] as well. At scale, TCloud can be integrated with a cluster-wise VM/container orchestration framework such as KubeVirt/Kubernetes [18] to decide which VM should be placed on what machine. Developing advanced vNPU/VM collocation policies is orthogonal to our work.
vNPU security isolation. TCloud enforces memory address space isolation among collocated vNPUs with the conventional memory segmentation scheme [57, 2] for both HBM and SRAM. TCloud divides the SRAM and HBM into fixed-sized segments and maps each segment to the virtual address space of a vNPU. For the NPU core in Table II, an SRAM/HBM segment is 2MB/1GB. There is no external fragmentation since the segment size is fixed. The address translation is performed by adding the segment offset to the starting address of the physical segment, which incurs negligible overhead. A page fault will be triggered when an invalid access happens. This is sufficient since ML frameworks like TensorFlow typically request a contiguous chunk of memory for the entire lifetime of an ML inference service and have their own memory management mechanism. To isolate the vNPU instances as they communicate with the host, TCloud uses IOMMU to enforce DMA remapping (§$\S$§III-F). We leave side-channel mitigation to future work.
@@ III-E Architectural Support for NeuISA
Next, the operation scheduler selects which operations from the instruction queues will be executed at every cycle. The ME operations from the ME $\mu$TOp instruction queues are directly issued to the corresponding MEs. For the VE operations, the scheduler selects which operations to issue from all VE $\mu$TOp instruction queues. To reclaim a harvested ME, TCloud performs a context switch to preempt the harvesting $\mu$TOp. Upon a context switch, the register file and the intermediate data in the MEs are saved to SRAM, which incurs negligible overhead compared to the length of an operator. The number of instruction queues should be large enough to support simultaneous execution of all MEs/VEs. For an NPU core with $n_{x}$ MEs and $n_{y}$ VEs, there are $n_{x}$ ME $\mu$TOp instruction queues and $n_{y}$ VE $\mu$TOp instruction queues.
@@ III-F System Support for NPU Virtualization
OS hypervisor. TCloud can work with OS hypervisors to provide system support for virtualizing NPUs. Take the KVM hypervisor as a case study, TCloud leverages vfio-mdev to expose vNPUs to VMs as mediated PCIe devices [30]. The hypervisor only mediates the resource management functions that are not on the critical path, including the following hypercalls: (1) create a new vNPU, (2) change the configuration of an existing vNPU, and (3) deallocate a vNPU. The hypercalls are routed to the vNPU manager, which is implemented as a host kernel module. The vNPU manager tracks the allocated and free resources (e.g., MEs/VEs, SRAM, HBM) of all physical NPUs on the host machine and implements the vNPU mapping policies (§III-C). Once a vNPU is set up, the VM can bypass the hypervisor and directly talk to the NPU device. TCloud uses SR-IOV [55] to expose each vNPU as a PCIe virtual function to the VM via PCIe-passthrough. The IOMMU performs DMA and interrupt remapping for the vNPUs.
Guest VM software. TCloud requires minimal changes to the guest VM software stack. First, the user source code remains unchanged. Typically, user codes are programmed with ML frameworks like PyTorch or TensorFlow [47, 22]. Second, for ML frameworks, only the backend NPU compiler needs to be revised to support NeuISA (§III-D). The ML framework has two parts: (1) The frontend converts the user code into a device-agnostic DNN dataflow graph and optionally partitions the graph onto multiple NPU cores. As our vNPU abstraction reflects the hierarchy of a physical NPU device, the frontend requires no changes. (2) The backend compiles the DNN graph into NPU binary using NeuISA. Third, the NPU vendor will provide a para-virtualized vNPU driver, which is a common practice for virtualizing PCIe devices. The vNPU driver provides user APIs for vNPU management and issues hypercalls to realize these APIs. With PCIe passthrough, the vNPU driver can directly interact with the NPU device [30].
@@
III-G TCloud Implementation
We implement TCloud with a production-level event-driven NPU simulator. We obtain the operator execution traces for each DNN workload on real Google Cloud TPUs. For each operator, the trace contains the ME/VE time, HBM time, tensor shapes, and the tile sizes and tiling dimensions selected by the compiler. We use the tiling information to generate $\mu$TOps and replay the generated $\mu$TOp traces in our simulator. We modify the frontend of the NPU simulator to implement the $\mu$TOp scheduling and harvesting policy (see §III-E). The scheduler picks $\mu$TOps from multiple traces (each trace represents the DNN workload of a vNPU) and issues them to the backend, which simulates the execution of each ME/VE, on-chip SRAM accesses, and DMA operations to the off-chip HBM at cycle level. To model the penalty of $\mu$TOp preemption (i.e., the context switch overhead of MEs), we set the ME preemption latency to 256 cycles based on the systolic array dimension (i.e., 128$\times$128), including 128 cycles to pop the partial sums and 128 cycles to pop the weights of the preempted $\mu$TOp.
We also prototype the hardware scheduler for NeuISA in Verilog and synthesize it using the FreePDK-15nm cell library [1]. Since the DNN workloads have deterministic dataflow graphs, they do not require complex dependency tracking or speculation in hardware. The hardware area overhead of TCloud is only 0.04% on a TPUv4 chip. The power overhead of this small extra area is negligible compared to that of the entire chip.

### v2

@@ Abstract
We present Neu10, a holistic NPU virtualization framework. We investigate virtualization techniques for NPUs across the entire software and hardware stack. Neu10 consists of (1) a flexible NPU abstraction called vNPU, which enables fine-grained virtualization of the heterogeneous compute units in a physical NPU (pNPU); (2) a vNPU resource allocator that enables pay-as-you-go computing model and flexible vNPU-to-pNPU mappings for improved resource utilization and cost-effectiveness; (3) an ISA extension of modern NPU architecture for facilitating fine-grained tensor operator scheduling for multiple vNPUs. We implement Neu10 based on a production-level NPU simulator. Our experiments show that Neu10 improves the throughput of ML inference services by up to 1.4$\times$ and reduces the tail latency by up to 4.6$\times$, while improving the NPU utilization by 1.2$\times$ on average, compared to state-of-the-art NPU sharing approaches.
Index Terms:
virtualization, neural processing unit, machine learning accelerator.
@@ I Introduction
Ideally, we wish to virtualize NPUs to enable flexible and fine-grained resource sharing and scheduling for improved NPU utilization and performance isolation. We present Neu10, a hardware-assisted system virtualization framework for NPUs.
@@
Neu10 can map vNPUs to physical compute units of NPU cores in different manners, based on the service level objectives (SLOs) of ML services. To maximize the NPU utilization while ensuring performance isolation, Neu10 enables fine-grained spatial sharing with resource harvesting. It also enables the oversubscription of NPU cores by temporally sharing MEs/VEs among multiple vNPUs. Therefore, the idle compute units can be opportunistically utilized by collocated workloads.
To facilitate the dynamic scheduling for collocated vNPUs, Neu10 extends the VLIW-style ISA by reorganizing VLIW instructions into independent micro-Tensor operators ($\mu$TOps in §$\S$§III). Neu10 introduces necessary architectural logic for fine-grained dynamic scheduling of $\mu$TOps on the shared physical NPU cores. It allows one vNPU to harvest available compute cycles of MEs/VEs from collocated vNPUs, without causing much interference. This is impossible with conventional VLIW-style ISAs, as they strictly couple the control flows of the (statically) allocated compute units. Our new architectural support enables Neu10 to offer the flexibility of NPU resource allocation and scheduling across the software (i.e., vNPU abstraction) and hardware (i.e., fine-grained $\mu$TOp scheduling) stack. Neu10 requires minimum modifications to NPU chips (0.04% die area cost) as well as ML compilers.
@@
We implement Neu10 with a production-level NPU simulator following the typical TPU architecture. We collect the traces of ML services as we run the MLPerf benchmarks [48] and the TPU reference models [23] on the real Google TPUs. Our experiments with multi-tenant ML instances show that Neu10 can improve the throughput of ML inference services by up to 1.4$\times$ and reduce the tail latency by up to 4.6$\times$, while improving the NPU utilization by 1.2$\times$ on average, in comparison with state-of-the-art NPU sharing approaches. We summarize the contributions of Neu10 as follows:
@@ III Design and Implementation
We design Neu10 to achieve the following objectives:
@@
vNPU abstraction. A vNPU instance reflects the hierarchy of a physical NPU board. Figure 10 shows the configurable parameters of a vNPU. Each vNPU is exposed to the VM as a PCIe device. The guest NPU driver can query the hierarchy of the vNPU, such as the number of chips, cores per chip, HBM size, and others. The maximum vNPU size is capped by the physical NPU size. If a guest VM requires more resources than is available on a physical NPU board, Neu10 can allocate multiple vNPU instances to it. The guest ML framework can handle the data distribution across multiple vNPU cores in the same way as that on physical NPUs. Take Google TPU for example, TensorFlow already handles data parallelism across physical NPUs. It can work in the same way with vNPUs.
Figure 11: System architecture of Neu10.
vNPU lifecycle. To create a vNPU instance, a user can specify the vNPU configuration following the pay-as-you-go model [50]. Cloud providers can define various default configurations (e.g., small/medium/large vNPU cores as having 1/4/8 MEs/VEs). Neu10 can also learn an optimized vNPU configuration for a DNN workload with ML compilers (§$\S$§III-B). As shown in Figure 11, upon vNPU initialization, the guest driver sends a request to the hypervisor through a para-virtualized interface (§III-F) ( 1). The vNPU manager maps the vNPU instance to NPU hardware resources (§$\S$§III-C). Then, it initializes the vNPU context in the physical NPU device and creates the MMIO mappings for the guest VM to access the vNPU ( 2). During execution, the application issues commands such as memcpy and compute offloading through the command buffer. The NPU hardware directly fetches the commands from the host memory without the hypervisor intervention. It also has DMA access to the DMA buffer in the guest memory space via the IOMMU. The DNN program on the NPU executes asynchronously from the CPU program, and the NPU hardware schedules vNPUs (§$\S$§III-E) independently of existing OS/hypervisor schedulers. The guest VM waits for the completion interrupt or actively polls the memory-mapped control registers for the current status of the vNPU ( 3). After execution, the user can free the vNPU.
@@
Following the popular pay-as-you-go model [50], cloud platforms allow users to specify the vNPU configuration on demand. However, as ML inference workloads have diverse ME/VE demands (see §$\S$§II-B), specifying the number of MEs/VEs can be challenging for users who are not NPU experts. Thus, we allow them to specify the total number of execution units (EUs), which is directly related to the cost of running the vNPU instance. Neu10 provides the vNPU allocator, a compile-time tool to improve the performance per cost of vNPUs by identifying an optimized ME/VE ratio for the user workload.
@@
Memory allocation. Users can use the compiler to estimate the total HBM capacity needed by a DNN workload. By default, the SRAM capacity is allocated proportionally to the number of allocated MEs, as more MEs usually indicate larger tile sizes. Based on our study in §$\S$§II-B, for many common ML inference services, the HBM bandwidth is less of a concern. Thus, Neu10 allows fair sharing of HBM bandwidth by default. For large models that demand large HBM capacity and bandwidth, the vNPU abstraction offers the flexibility for end users to allocate the demanded resources. The user may also leverage existing tensor swapping techniques to support large DNN workloads with limited memory capacity [28, 63]. After vNPU allocation, ML compilers will compile the DNN program with the allocated resources. The compiler ensures the DNN program does not exceed the allocated SRAM and HBM. We will discuss how Neu10 handles compilation for different numbers of MEs/VEs in §$\S$§III-D.
@@
The vNPU manager attempts to balance the number of allocated EUs and the size of allocated memory. This minimizes the chance that all EUs on one core are allocated but a large portion of its memory is not allocated, or vice versa. Thus, vNPUs with many EUs and small memory will be collocated with vNPUs with few EUs and large memory. Neu10 uses a greedy algorithm for this by default.
vNPU mapping schemes. Neu10 provides the flexibility for cloud platforms to enable both hardware-isolated (spatial-isolated) mapping and software-isolated (temporal-sharing) mapping. With hardware-isolated mapping, a vNPU is mapped to dedicated EUs and SRAM, and the allocated hardware is not shared with other vNPUs. With software-isolated mapping, multiple vNPUs can temporally share the same EUs. Neu10 uses priority-based scheduling for fair sharing and performs context switches between vNPUs (see §$\S$§III-E).
vNPU mapping policies. Neu10 decides which vNPUs can be mapped onto the same physical NPU (pNPU) as follows. With hardware-isolated mapping, Neu10 collocates a set of vNPUs as long as the total resource requirement (e.g., number of MEs/VEs, HBM capacity) does not exceed the pNPU. With software-isolated mapping, Neu10 aims to load-balance the pNPUs while allowing oversubscription. Neu10 tracks the total resource requirement of assigned vNPUs on each pNPU, and assigns a new vNPU to the pNPU that suffers the least resource requirement. Neu10 can support other collocation policies [41, 57, 12] as well. At scale, Neu10 can be integrated with a cluster-wise VM/container orchestration framework such as KubeVirt/Kubernetes [18] to decide which VM should be placed on what machine. Developing advanced vNPU/VM collocation policies is orthogonal to our work.
vNPU security isolation. Neu10 enforces memory address space isolation among collocated vNPUs with the conventional memory segmentation scheme [57, 2] for both HBM and SRAM. Neu10 divides the SRAM and HBM into fixed-sized segments and maps each segment to the virtual address space of a vNPU. For the NPU core in Table II, an SRAM/HBM segment is 2MB/1GB. There is no external fragmentation since the segment size is fixed. The address translation is performed by adding the segment offset to the starting address of the physical segment, which incurs negligible overhead. A page fault will be triggered when an invalid access happens. This is sufficient since ML frameworks like TensorFlow typically request a contiguous chunk of memory for the entire lifetime of an ML inference service and have their own memory management mechanism. To isolate the vNPU instances as they communicate with the host, Neu10 uses IOMMU to enforce DMA remapping (§$\S$§III-F). We leave side-channel mitigation to future work.
@@ III-E Architectural Support for NeuISA
Next, the operation scheduler selects which operations from the instruction queues will be executed at every cycle. The ME operations from the ME $\mu$TOp instruction queues are directly issued to the corresponding MEs. For the VE operations, the scheduler selects which operations to issue from all VE $\mu$TOp instruction queues. To reclaim a harvested ME, Neu10 performs a context switch to preempt the harvesting $\mu$TOp. Upon a context switch, the register file and the intermediate data in the MEs are saved to SRAM, which incurs negligible overhead compared to the length of an operator. The number of instruction queues should be large enough to support simultaneous execution of all MEs/VEs. For an NPU core with $n_{x}$ MEs and $n_{y}$ VEs, there are $n_{x}$ ME $\mu$TOp instruction queues and $n_{y}$ VE $\mu$TOp instruction queues.
@@ III-F System Support for NPU Virtualization
OS hypervisor. Neu10 can work with OS hypervisors to provide system support for virtualizing NPUs. Take the KVM hypervisor as a case study, Neu10 leverages vfio-mdev to expose vNPUs to VMs as mediated PCIe devices [30]. The hypervisor only mediates the resource management functions that are not on the critical path, including the following hypercalls: (1) create a new vNPU, (2) change the configuration of an existing vNPU, and (3) deallocate a vNPU. The hypercalls are routed to the vNPU manager, which is implemented as a host kernel module. The vNPU manager tracks the allocated and free resources (e.g., MEs/VEs, SRAM, HBM) of all physical NPUs on the host machine and implements the vNPU mapping policies (§III-C). Once a vNPU is set up, the VM can bypass the hypervisor and directly talk to the NPU device. Neu10 uses SR-IOV [55] to expose each vNPU as a PCIe virtual function to the VM via PCIe-passthrough. The IOMMU performs DMA and interrupt remapping for the vNPUs.
Guest VM software. Neu10 requires minimal changes to the guest VM software stack. First, the user source code remains unchanged. Typically, user codes are programmed with ML frameworks like PyTorch or TensorFlow [47, 22]. Second, for ML frameworks, only the backend NPU compiler needs to be revised to support NeuISA (§III-D). The ML framework has two parts: (1) The frontend converts the user code into a device-agnostic DNN dataflow graph and optionally partitions the graph onto multiple NPU cores. As our vNPU abstraction reflects the hierarchy of a physical NPU device, the frontend requires no changes. (2) The backend compiles the DNN graph into NPU binary using NeuISA. Third, the NPU vendor will provide a para-virtualized vNPU driver, which is a common practice for virtualizing PCIe devices. The vNPU driver provides user APIs for vNPU management and issues hypercalls to realize these APIs. With PCIe passthrough, the vNPU driver can directly interact with the NPU device [30].
@@
III-G Neu10 Implementation
We implement Neu10 with a production-level event-driven NPU simulator. We obtain the operator execution traces for each DNN workload on real Google Cloud TPUs. For each operator, the trace contains the ME/VE time, HBM time, tensor shapes, and the tile sizes and tiling dimensions selected by the compiler. We use the tiling information to generate $\mu$TOps and replay the generated $\mu$TOp traces in our simulator. We modify the frontend of the NPU simulator to implement the $\mu$TOp scheduling and harvesting policy (see §III-E). The scheduler picks $\mu$TOps from multiple traces (each trace represents the DNN workload of a vNPU) and issues them to the backend, which simulates the execution of each ME/VE, on-chip SRAM accesses, and DMA operations to the off-chip HBM at cycle level. To model the penalty of $\mu$TOp preemption (i.e., the context switch overhead of MEs), we set the ME preemption latency to 256 cycles based on the systolic array dimension (i.e., 128$\times$128), including 128 cycles to pop the partial sums and 128 cycles to pop the weights of the preempted $\mu$TOp.
We also prototype the hardware scheduler for NeuISA in Verilog and synthesize it using the FreePDK-15nm cell library [1]. Since the DNN workloads have deterministic dataflow graphs, they do not require complex dependency tracking or speculation in hardware. The hardware area overhead of Neu10 is only 0.04% on a TPUv4 chip. The power overhead of this small extra area is negligible compared to that of the entire chip.